A 45 n toolbox is dragged horizontally at constant speed by a rope making an angle of 40 with the floor the tension in the rope is 20 n, what is the coefficient of kinetic friction?
1
Expert's answer
2019-06-17T12:59:36-0400

The freebody diagram of the considered system is shown below.



The net force equals 0 because the speed is constant. Hence, we obtain:


"\vec{N}+\vec{T}+\vec{mg} +\vec{F}_{fr} =0"

which can be written in terms of x- and y-projections as:


"F_{fr} = T \cos{\alpha}\\
N + T \sin{\alpha} = mg"

Taking into account that


"F_{fr} = \mu N"

one can derive:


"\mu = \frac{T \cos{\alpha} }{N} = \frac{T \cos{\alpha}}{mg - T \sin{\alpha}}"

Substituting the numerical values, we obtain:


"\mu = \frac{20 \cos{40^\circ}}{40 - 20 \sin{40 ^\circ}} \approx 0.56"

Answer: 0.56